# Large Language Models for Automated Literature Review: An Evaluation of Reference Generation, Abstract Writing, and Review Composition

### v3

@@ Abstract
Large language models (LLMs) have emerged as a potential solution to automate the complex processes involved in writing literature reviews, such as literature collection, organization, and summarization. However, it is yet unclear how good LLMs are at automating comprehensive and reliable literature reviews. This study introduces a framework to automatically evaluate the performance of LLMs in three key tasks of literature writing: reference generation, literature summary, and literature review composition. We introduce multidimensional evaluation metrics that assess the hallucination rates in generated references and measure the semantic coverage and factual consistency of the literature summaries and compositions against human-written counterparts. The experimental results reveal that even the most advanced models still generate hallucinated references, despite recent progress. Moreover, we observe that the performance of different models varies across disciplines when it comes to writing literature reviews. These findings highlight the need for further research and development to improve the reliability of LLMs in automating academic literature reviews.
@@ 1 Introduction
The literature review is a critical component of academic writing that aims to synthesize, critique, and assess the current state of knowledge in a particular field. It involves a comprehensive examination of published research articles, theoretical frameworks, and research methodologies related to a specific topic. Conducting a thorough literature review often necessitates extensive reading and summarizing of pertinent literature, which can be a complex and time-consuming process, especially in well-established fields where the number of relevant references can range from dozens to hundreds. To alleviate this burden, researchers have recently turned to advanced deep learning models as a potential tool to aid in the automated generation of literature reviews. Kontonatsios et al. (2020); Aliyu et al. (2018). The emergence of large language models (LLMs) has introduced a promising avenue for automating various aspects of the literature review writing process. LLMs, trained on extensive collections of academic texts, have developed impressive generative capabilities that make them a potent tool for assisting researchers in crafting literature reviews. By leveraging the knowledge acquired from vast amounts of scholarly literature, these models can help researchers quickly identify relevant sources, summarize key findings, and generate coherent syntheses of the existing body of work Wang et al. (2024b); Agarwal et al. (2024).
The process of writing a literature review requires a deep understanding of the studies within a particular field, as well as the ability to summarize the core contributions of the relevant literature. While techniques such as Retrieval-Augmented Generation (RAG) can enhance the domain-specific knowledge of LLMs, most researchers rely on vanilla LLMs, such as ChatGPT, for literature review writing without the use of RAG Wang et al. (2024a). Consequently, it is crucial to evaluate the performance of these native LLMs in the context of literature review writing to determine their effectiveness and limitations.
Therefore, in this paper, we propose a framework for automatically assessing the literature review writing ability of LLMs, using human-written literature reviews as the gold standard and designing metrics for a comprehensive evaluation. We first collect a dataset of human-written literature reviews to serve as a benchmark for evaluating the performance of LLMs. We then ask LLMs to complete three tasks based on the collected dataset: generating references, writing an abstract, and writing a complete literature review based on a given topic. Finally, we evaluate the generated results from several dimensions, including the presence of hallucinations in the references, as well as the semantic coverage and factual consistency of the generated literature review compared to the human-written context. By assessing the performance of LLMs across these tasks and evaluating their output using our proposed metrics, we aim to provide a comprehensive understanding of their capabilities and limitations in writing literature reviews.
@@
  • Second, we collect 1,105 literature reviews from 51 journals across 5 disciplines as the ground truth. We then design three tasks for accessing LLMs in literature writing: reference generation, literature summary, and literature composition on a given topic.
@@ 2 Related Work
Recent studies have increasingly focused on using LLMs for writing literature reviews. For example,  Wang et al. (2024b) proposed a literature survey writing process for LLMs, called AutoSurvey, which incorporates real-time knowledge by retrieving the latest relevant papers using a retrieval-augmented generation (RAG) approach. Similarly,  Agarwal et al. (2024) investigated the zero-shot capabilities of recent LLMs in generating literature reviews from abstracts. They employed a two-step approach: first retrieving relevant papers via an external knowledge base, then outlining and executing a structured plan to generate the review. The authors focused on the LLMs’ ability to effectively summarize the literature.
Additionally, for evaluating the literature review writing capabilities of LLMs, more attention has been paid to assessing hallucinations in reference citations. For example, Chelli_Descamps_Lavoué_Trojani_Azar_Deckert_Raynier_Clowez_Boileau_Ruetsch-Chelli_2024 used LLMs (ChatGPT, GPT-4, and Bard LLM) to generate 11 systematic reviews on shoulder rotator cuff pathology and evaluated the hallucination rates of the references. They found that Bard had significantly higher rates of hallucinations than ChatGPT and GPT-4. Similarly, Agrawal et al. (2024) chose 200 computer science topics to evaluate LLMs’ hallucination rates. They first asked LLMs to generate 5 reference titles, then used the BING Search API to automatically label the references as true or not. The LLMs were then examined to determine if they could actively discover the hallucinated literature through direct queries (asking directly if the paper exists) and indirect queries (who wrote the paper?). Athaluri et al. (2023) assessed the hallucinations of LLMs in research proposal writing. They used ChatGPT to generate 50 research proposals based on 50 given topics. ChatGPT generated 178 references in these proposals, and then they employed five researchers to manually check these references and DOIs on Scopus, Google, and PubMed search engines. They found that among the references, 109 had valid DOIs. Moreover, Aljamaan et al. (2024) provided 5 medical topics and used ChatGPT 3.5, Bard, Perplexity, Bing, Elicit, and SciSpace to generate 10 relevant references for each topic, with each reference including 7 items: title, journal name, author, DOI, publication date, link to the paper, and relevance of the reference to the keyword prompt. They then proposed the Reference Hallucination Score (RHS) for citation hallucination assessment of LLMs, assigning different hallucination scores to each item, e.g., a score of 2 corresponding to the title and 1 corresponding to the publication date. While these studies are very useful attempts at evaluating citation hallucinations in literature reviews, they are limited to specific academic topics and do not provide a comprehensive assessment of LLM performance. Moreover, these evaluations often involve manual human intervention.
@@ 3 Methodology
In this section, we propose a framework for evaluating LLMs’ literature survey writing ability. The framework, as shown in Figure 1, consists of three main stages: dataset construction and task design for evaluation, collection LLM-generated output, and assessment of the generated output.
@@ 3.1 Dataset Construction
Then, the dataset $D$ is the article set from 51 journals, $D=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles. Each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{k},...,r_{K}\}$, and $K$ represents the length of the reference set.
@@ 3.2 Task Design
Since literature review writing primarily involves the collection and synthesis of relevant research, we design three tasks as follows to evaluate LLMs’ capabilities in different aspects of literature review writing.
  • Task 1: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ to the research topic. Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page. In this task, we evaluate whether LLMs can recommend reliable references based on the given topic.
  • Task 2: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to write an abstract $a_{i}^{g}$according to the research topic provided in the title and the keywords. The length of the generated abstract is to be the same as the length of the original abstract of the article. In this task, we evaluate whether LLMs can write an abstract based on the given topic.
  • Task 3: Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a short literature review $c_{i}^{g}$ according to the research topic provided in the title, keywords, and abstract. The literature survey should be about 1000 words long. LLMs also need to back up claims by citing previous studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ (with a total of $N$ citations in the literature survey). In this task, we evaluate whether LLMs can write a high-quality literature survey and cite truth studies.
@@ 3.3 Evaluation Metrics
Reference hallucination evaluation metrics. Given that LLMs are trained on vast corpora, including academic sources, we aim to evaluate whether they can generate true references. In this section, we introduce the calculation process of the reference accuracy $Acc.$ and title search rate $S_{t}$ for each LLM. A higher accuracy metric indicates a lower hallucination rate.
@@
For each article $p_{i}\in D$, each LLM generates $N$ references $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ in Task 1 and Task 3, each $r_{n}^{g}=\{T,A,J,Y,V,FP,LP\}$ and includes 7 elements corresponding to the generated article title, authors, journal, year, volume, first page, and last page. Each element corresponding to a state label represents whether it is accurate or not $\{e_{d}\mid d=0,1,\dots,7\}$, $e_{d}$ = 1 or 0.
@@
Next, we describe how to obtain $\{e_{d}\mid d=0,1,\dots,7\}$. First, we use the generated titles $T$ and the first author in $A$ as the queries and search them separately from external academic search engines. Then, we obtain two candidate sets $Z_{t}$ and $Z_{a}$ respectively. Subsequently, we compare the generated $r_{n}^{g}$ with the article from candidate sets $Z_{t}$ and $Z_{a}$. For example, if the title of the article candidate is the same as that of $r_{n}^{g}$, then $e_{0}=1$. Finally, we find the best candidate article based on the sum of $\{e_{d}\mid d=0,1,\dots,6\}$, and the one with the largest sum is the best candidate article $z_{n}^{g}$ of $r_{n}^{g}$.
Then, we compare the algin degree between the generated $r_{n}^{g}$ and the best candidate article $z_{n}^{g}$ to determine whether $r_{n}^{g}$ corresponds to a true article (As shown in Eq 1). On the one hand, when its title $T$ is correct and one other element is correct, we consider $r_{n}^{g}$ corresponds to a true article. Here its title $T$ is correct or not, we give LLM a certain margin of error, as long as the title has a match rate of 80%, it will be judged as accurate, so we need another piece of information that is correct to ensure that the article is reliable. On the other hand, when its title $T$ is incorrect, certain other information is correct which can also help us to locate the article. Therefore, when the other three elements in $r_{n}^{g}$ are matching, we also consider it as a reliable article.
@@
$\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}$ (1)
@@
For each paper $p_{i}$ in the dataset, we formulate the accuracy of the references generated by the LLM as Eq 2. Then, we obtain the accuracy of generated references for each LLM $Acc.$ by averaging $Acc_{ref.}(p_{i})$ across articles, as shown in Eq 3.
$Acc_{ref.}(p_{i})=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})$ (2)
$Acc.=\frac{1}{M}\sum_{i=0}^{M}Acc_{ref.}(p_{i})$ (3)
On the other hand, intuitively, the title is the key element that reflects the generated reference’s faithfulness. In  Agrawal et al. (2024)’s work, they determined by ground-truth labels assigned using the Bing search API. Encouraged by this work, we also calculate the title search scores of generated references for each LLM, formulated as follows.
$S_{t}=\frac{1}{MN}\sum_{M}\sum_{N}s_{p_{i}}$ (4)
@@
$s_{p_{i}}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from external Scholar API,}\\ 0&\text{otherwise}\end{cases}$ (5)
where $s_{p_{i}}$ is the number of the generated title that can be matched in the external scholar base. We search the generated title from an external academic search engine to estimate whether the title can be matched.
@@
Context evaluation metrics. In this paper, we use the human-written article as the gold truth and then evaluate LLM-generated context from factual consistency and semantic coverage aspects. The resemblance of natural language inference (NLI) to factual consistency evaluation has led to utilizing NLI models for measuring factual consistency Gao et al. (2023). Encouraged by previous works, we also use the NLI method to evaluate the factual consistency between LLMs generated and human-written text. For example, we calculate the NLI score $Entail_{p_{i}}$ between the original article abstract $a_{i}$ and the LLM-generated abstract $a_{i}^{g}$ as follows.
@@
$Entail=\frac{1}{M}\sum_{i=0}^{M}{Entail_{p_{i}}}$ (7)
Additionally, we use commonly employed semantic similarity metrics and Key Point Recall (KPR) to calculate the semantic coverage between the context generated by LLMs and human-written context. Specifically, for Task 2, we apply cosine similarity and the ROUGE metric for semantic coverage evaluation. For Task 3, we use the ROUGE metric and KPR to measure the semantic coverage of the literature review generated by the LLMs relative to human-written content.
@@
$KPR=\frac{1}{M}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}$ (8)

### v4

@@ Abstract
Large language models (LLMs) have emerged as a potential solution to automate the complex processes involved in writing literature reviews, such as literature collection, organization, and summarization. However, it is yet unclear how good LLMs are at automating comprehensive and reliable literature reviews. This study introduces a framework to automatically evaluate the performance of LLMs in three key tasks of literature review writing: reference generation, abstract writing, and literature review composition. We introduce multidimensional evaluation metrics that assess the hallucination rates in generated references and measure the semantic coverage and factual consistency of the literature summaries and compositions against human-written counterparts. The experimental results reveal that even the most advanced models still generate hallucinated references, despite recent progress. Moreover, we observe that the performance of different models varies across disciplines when it comes to writing literature reviews. These findings highlight the need for further research and development to improve the reliability of LLMs in automating academic literature reviews. The dataset and code used in this study are publicly available at an anonymous repository  [^1] https://anonymous.4open.science/r/Eval_LLM_LR-C657.
Large Language Models for Automated Literature Review: An Evaluation of Reference Generation, Abstract Writing, and Review Composition
Xuemei Tang1 Xufeng Duan1 Zhenguang G. Cai1,2 1Department of Linguistics and Modern Languages, The Chinese University of Hong Kong 2Brain and Mind Institute, The Chinese University of Hong Kong
@@ 1 Introduction
The literature review is a critical component of academic writing that aims to synthesize, critique, and assess the current state of knowledge in a particular field. It involves a comprehensive examination of published research articles, theoretical frameworks, and research methodologies related to a specific topic. Conducting a thorough literature review often necessitates extensive reading and summarizing of pertinent literature, which can be a complex and time-consuming process, especially in well-established fields where the number of relevant references can range from dozens to hundreds. To alleviate this burden, researchers have recently turned to advanced deep learning models as a potential tool to aid in the automated generation of literature reviews Aliyu et al. (2018); Kontonatsios et al. (2020).
The emergence of large language models (LLMs) has introduced a promising avenue for automating key aspects of literature review writing, including identifying relevant sources, summarizing findings, and generating coherent syntheses Wang et al. (2024b); Agarwal et al. (2024); Hsu et al. (2024). While techniques such as Retrieval-Augmented Generation (RAG) can enhance the domain-specific knowledge of LLMs, most researchers rely on vanilla LLMs, such as ChatGPT, for literature review writing without the use of RAG Wang et al. (2024a). Consequently, it is crucial to evaluate the performance of these naive LLMs in the context of literature review writing to determine their effectiveness and limitations.
Therefore, in this paper, we propose a framework for automatically assessing the literature review writing ability of LLMs, using human-written literature reviews as the gold standard and designing metrics for a comprehensive evaluation. We first collect a dataset of human-written literature reviews to serve as a benchmark for evaluating the performance of LLMs. We then ask LLMs to complete three tasks based on the collected dataset: generating references, writing an abstract, and writing a complete literature review based on a given topic. Finally, we evaluate the generated results from several dimensions, including the presence of hallucinations in the references, as well as the semantic coverage and factual consistency of the generated abstract and literature review compared to the human-written context. By assessing the performance of LLMs across these tasks and evaluating their output using our proposed metrics, we aim to provide a comprehensive understanding of their capabilities and limitations in writing literature reviews.
@@
  • Second, we collect 1,105 literature reviews from 51 journals across 5 disciplines as the ground truth. We then design three tasks for accessing LLMs in literature writing: reference generation, abstract writing, and literature composition on a given topic.
@@ 2 Related Work
Recent studies have explored LLMs for literature review generation. For example, Wang et al. (2024b) proposed AutoSurvey, which incorporates up-to-date papers via retrieval-augmented generation. Agarwal et al. (2024) examined zero-shot LLM review generation using a two-step retrieval and outlining process. More recently, Liang et al. (2025) presented SurveyX, an efficient system that optimizes retrieval, extraction, and outline generation, supporting multimodal outputs such as figures and tables.
Additionally, recent efforts to evaluate literature review generation by LLMs have increasingly focused on assessing hallucinations in reference citations. For instance, Chelli et al. (2024) analyzed hallucination rates in 11 systematic reviews on shoulder rotator cuff pathology generated by ChatGPT, GPT-4, and Bard, finding Bard exhibited significantly higher hallucination rates. Similarly, Agrawal et al. (2024) evaluated hallucinations across 200 computer science topics by generating reference titles with LLMs and verifying their existence via the BING Search API, further probing whether LLMs could detect hallucinated references through direct and indirect queries. Athaluri et al. (2023) examined hallucinations in 50 ChatGPT-generated research proposals, manually validating references and DOIs using Scopus, Google, and PubMed, reporting 109 valid DOIs among 178 references. Additionally, Aljamaan et al. (2024) introduced the Reference Hallucination Score (RHS) by generating references for five medical topics across multiple LLMs, assigning weighted hallucination scores to citation components such as title and publication date. While these studies provide valuable insights, they are generally limited to specific domains and rely heavily on manual evaluation, lacking a comprehensive, scalable assessment framework for LLM reference hallucinations.
@@ 3 Methodology
In this section, we propose a framework for evaluating LLMs’ literature review writing ability. The framework, as shown in Figure 1, consists of three main stages: dataset construction and task design for evaluation, collection LLM-generated output, and assessment of the generated output.
Figure 1: Illustration of the evaluation framework.
@@ 3.1 Dataset Construction
Then, the dataset $D$ is the article set from 51 journals, $D=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles. Each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r_{1},...,r_{k},...,r_{K}\}$, and $K$ represents the length of the reference set.
@@ 3.2 Task Design
Since literature review writing primarily involves the collection and synthesis of relevant research, we design three independent tasks as follows to evaluate LLMs’ capabilities in different aspects of literature review writing.
  • Reference Generation: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ to the research topic. Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page, $r_{n}^{g}=\{T,A,J,Y,V,FP,LP\}$. In this task, we evaluate whether LLMs can recommend reliable references based on the given topic. Note that these references are not reused in later tasks.
  • Abstract Writing: Given an article title $t_{i}$ and its associated keywords $w_{i}$, the LLMs are prompted to generate an abstract $a_{i}^{g}$ that aligns with the research topic. The length of the generated abstract is constrained to match that of the original. This task serves as a proxy for literature review planning, as abstracts often outline the key components of a study—such as its objectives, methods, and covered subtopics—which are also critical in structuring comprehensive literature reviews. By evaluating the model’s ability to generate coherent and topic-relevant abstracts, we assess its potential to assist researchers in the early planning stages of literature review writing.
  • Review Composition: Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a short literature review $c_{i}^{g}$ according to the research topic provided in the title, keywords, and abstract. To facilitate evaluation and accommodate computational budget constraints, the length of each literature review is limited to approximately 1000 words. LLMs also need to back up claims by citing relevent studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ (with a total of $N$ citations in the literature review). These citations are newly generated to support the content of the review. In this task, we evaluate whether LLMs can write a high-quality literature review and cite truth studies.
@@ 3.3 Evaluation Metrics
Reference hallucination evaluation metrics. Given that LLMs are trained on vast corpora, including academic sources, we aim to evaluate whether they can generate true references. In this section, we introduce the calculation process of the reference precision $Precison$, reference recall $Recall$, $F1$, and title search rate $S_{t}$ for each LLM. A higher precision metric indicates a lower hallucination rate. A higher recall indicates that the LLM-generated references cover more of the ground-truth citations used by human authors, reflecting a better ability to identify key prior work relevant to the topic.
@@
For each article $p_{i}\in D$, each LLM generates $N$ references $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ in both Reference Generation and Review Composition tasks, each $r_{n}^{g}$ and includes 7 elements. Each element corresponding to a state label represents whether it is accurate or not $\{e_{d}\}_{d=0}^{6}$, $e_{d}$ = 1 or 0.
@@
Next, we describe how to obtain $\{e_{d}\}_{d=0}^{6}$. First, we use the generated titles $T$ and the first author in $A$ as the queries and search them separately from external academic search engines. This results in two sets of candidate articles, $Z_{t}$ and $Z_{a}$ respectively, We then merge the two sets and remove duplicates to obtain the final candidate set $Z=\{z_{1},...,z_{j},...,z_{J}\}$. Subsequently, we compare the generated $r_{n}^{g}$ with the article $z_{j}$ from candidate sets $Z$. For example, if the title of a candidate article $z_{j}$ matches the title of $r_{n}^{g}$, then $e_{0}=1$. Finally, we find the best candidate article based on the sum of $\{e_{d}\}_{d=0}^{6}$, and the one with the largest sum is the best candidate article $z_{j}$ of $r_{n}^{g}$.
Then, we compare the alignment degree between the generated reference $r_{n}^{g}$ and the best-matching candidate article $z_{j}$ to determine whether $r_{n}^{g}$ corresponds to a real article (as shown in Eq. 1). We consider $r_{n}^{g}$ to be reliable under either of the following two conditions:
Title-based matching: If the title $T$ is correct (i.e., $e_{0}=1$), and at least one other metadata element (e.g., author, journal, year, etc.) also matches, the reference is deemed reliable. To allow for minor variations, we consider the title to be correct if it achieves a match rate of at least 80% with the ground-truth title—a threshold determined through human evaluation.
Metadata-based matching: If the title $T$ is incorrect (i.e., $e_{0}=0$), we still consider the reference reliable if at least three of the remaining metadata elements (author, journal, year, volume, first page, last page) match those of a real article. This allows us to identify true references even when the title is noisy or incomplete.
@@
$\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}\vspace{-2mm}$ (1)
@@
For each paper $p_{i}$ in the dataset, we compute the reference precision of the LLM-generated references, denoted as $\textit{Precision}{(p_{i})}$, as defined in Eq. 2. We then obtain the overall Precision score for each LLM by averaging $\textit{Precision}{(p_{i})}$ across all papers in the dataset, as shown in Eq. 3.
$Precision_{(p_{i})}=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})\vspace{-2mm}$ (2)
$Preicison=\frac{1}{M}\sum_{i=0}^{M}Precison{(p_{i})}\vspace{-1mm}$ (3)
Precision is measured by comparing the LLM-generated references with external academic databases. We also evaluate Recall by comparing the references generated by the LLM with those cited in the human-written original articles. The key difference between precision and recall in our setting lies in the candidate set $Z$: for precision, $Z$ is constructed from external academic search results, whereas for recall, $Z$ consists of the references actually cited in the human-written articles. Based on precision and recall, we further compute the F1 score, which serves as a harmonic mean to reflect the overall accuracy of the reference generation.
Additionally, the title is intuitively the most critical element in determining the faithfulness of a generated reference. In the work of Agrawal et al. (2024), ground-truth labels were assigned based on results returned by the Bing Search API. Inspired by their approach, we also calculate the title search score for each LLM to estimate how many generated titles correspond to real publications.
$S_{t}=\frac{1}{MN}\sum_{M}\sum_{N}s_{p_{i}}^{(n)}\vspace{-2mm}$ (4)
@@
$s_{p_{i}}^{(n)}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from external Scholar API,}\\ 0&\text{otherwise}\end{cases}\vspace{-1mm}$ (5)
Here, $s_{p_{i}}^{(n)}$ indicates whether the generated title in reference $r_{n}^{g}$ for paper $p_{i}$ can be found using an external academic search engine. This metric helps estimate the proportion of references with verifiable titles among the total generated references.
@@
Context evaluation metrics. In our study, we use the human-written article as the gold truth and then evaluate LLM-generated context from factual consistency and semantic coverage aspects. The resemblance of natural language inference (NLI) to factual consistency evaluation has led to utilizing NLI models for measuring factual consistency Gao et al. (2023). Encouraged by previous works, we also use the NLI method to evaluate the factual consistency between LLMs generated and human-written text. For example, we calculate the NLI score $Entail_{p_{i}}$ between the original article abstract $a_{i}$ and the LLM-generated abstract $a_{i}^{g}$ as follows.
@@
$Entail=\frac{1}{M}\sum_{i=0}^{M}{Entail_{p_{i}}}\vspace{-2mm}$ (7)
Additionally, we use commonly employed semantic similarity metrics and Key Point Recall (KPR) to calculate the semantic coverage between the context generated by LLMs and human-written context. Specifically, for the Abstract Writing task, we apply cosine similarity and the ROUGE metric for semantic coverage evaluation. For the Review Composition task, we use the ROUGE metric and KPR to measure the semantic coverage of the literature review generated by the LLMs relative to human-written content.
@@
$KPR=\frac{1}{M}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}\vspace{-2mm}$ (8)